GIANT TORSIONAL SPLITTINGS IN THE PARALLEL FUNDAMENTAL BAND (v_{5}=1←0) OF CH_{3}SiH_{3}

J. SCHRODERUS^{a}, Department of Physics and Astronomy, University of British Columbia, Vancouver, B.C., Canada, V6T 1Z1; N.MOAZZEN-AHMADI, University of Lethbridge, Lethbridge, Alberta, Canada, T1K 3M4; \underline{I. OZIER}, Department of Physics and Astronomy, University of British Columbia, Vancouver, B.C., Canada, V6T 1Z1.

The lowest frequency parallel fundamental band ν_{5} of CH_{3}SiH_{3} near 700 cm^{−1} has been measured at a resolution of 0.004 cm^{−1} with Fourier transform spectroscopy. The torsional fine structure in the spectrum is increased from ∼0.005 to ∼1 cm^{−1} by the anharmonic interactions between the torsional stack of levels (v_{6}=0,1,2..) in the ground vibrational state and the corresponding stack with (v_{5}=1). For given (v_{5}, v_{6}, J, k), the torsional sub-levels are labelled by the index σ=0,+1,–1. For several (k,σ), resonant or near-resonant interactions occur between the vibration-torsion levels (v_{5}=0, v_{6}=5) and (v_{5}=1, v_{6}=0). For (k=1, σ=–1), the mixing is so severe that perturbation-allowed transitions (v_{6}=5←0) have been observed. In spite of the fact that v_{6} changes by 5 units, the mixing causes the probability density as a function of the torsional angle to have a local maximum for (v_{5}=1, v_{6}=0) in the centre of the classically forbidden region inside the torsional barrier. This severe mixing with such large |Δv_{6}| has serious implications for vibration-rotation relaxation in molecules with low frequency, highly anharmonic vibrational modes. A detailed fit to over 3400 transition frequencies has been made to within experimental accuracy using a vibration-torsion-rotation Hamiltonian with 45 parameters. The treatment includes a third stack of torsional levels in the lowest-lying degenerate mode ν_{12} near 525 cm^{−1}. The Hamiltonian is severely constrained by a wide variety of infrared, microwave and molecular beam transitions. The analysis shows that the second term in the Fourier expansion of the hindering potential characterized by the parameter V_{6} is dominantly due to these Fermi-type interactions. The changes in the effective molecular parameters caused by removing the non-resonant anharmonic interactions by contact transformation will be discussed.

a Present address: Department of Physical Sciences, University of Oulu, P.O. Box 3000, Oulu, Finland, FIN 90401